# Deep Transfer Learning: Model Framework and Error Analysis

### v1

@@ 2.1.2 Loss for representation learning
Notably, for any bijective mapping $\psi$ acting on $\mathbf{h}^{*}$, the composition $\psi\circ\mathbf{h}^{*}$ preserves the sufficiency and invariance properties required in (1)-(2). Therefore, we can constrain the distribution of $\mathbf{h}^{*}(\mathbf{X})$ to match the uniform distribution $\mathcal{U}_{r}$ without loss of generality, as the existence of such $\mathbf{h}^{*}(\mathbf{X})$ can be guaranteed by optimal transport theory [43]. Given that the entries of $\mathcal{U}_{r}$ are independent, we promote the disentanglement of $\mathbf{h}^{*}(\mathbf{X})$.
@@
$R(\mathbf{F},\mathbf{h})=$ $\mathbb{E}[\|Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S})\|^{2}]+\lambda\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ (10)
$+\tau\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1},$
where the first term is the regression loss for (3), and $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ is a measure used to evaluate independence between $\mathbf{h}(\mathbf{X})$ and $S$ as required in (2), such as the distance covariance (detail definition see Section 2.2.1, and
@@
$\hat{R}(\mathbf{F},$ $\mathbf{h})=\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}_{i}\mathbf{h}(X_{i,j}))^{2}+$ (11)
$\lambda\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]+\tau\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1},$
where $\sum_{j=1}^{p}n_{p}=n$, and $\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]$ is the empirical form of $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ (see Section 2.2.1 for detail), and
@@
denotes the empirical form of the Wasserstein distance, wherein $\{\xi_{i}\}_{i=1}^{n}$ denotes i.i.d random samples drawn from $\mathcal{U}_{r}$ and the function class $\mathcal{G}$ is chosen as the norm-constrained neural network class whose Lipschitz constants are bounded by $K_{3}$, see Section 2.2.2 for detail.
@@
$(\hat{\mathbf{F}},\hat{\mathbf{h}})=\mathop{\rm argmin}_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\hat{R}(\mathbf{F},\mathbf{h}),$ (12)
@@ 2.1.3 Loss for downstream task
$(\hat{\mathbf{F}}_{T},\hat{Q})\in\mathop{\rm argmin}_{(\mathbf{F}_{T},\mathbf{A},q)\in\mathcal{B}_{r}(B)\circ\mathcal{B}_{dd^{*}}(B)\circ\mathcal{Q}}\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ (13)
@@
2.2.1 Distance covariance
We recall the concept of distance covariance [44], which characterizes the dependence of two random variables. Let $\mathfrak{i}$ be the imaginary unit $(-1)^{1/2}$. For any $\mathbf{t}\in\mathbb{R}^{d}$ and $\mathbf{s}\in\mathbb{R}^{m}$, let $\psi_{Z}(\mathbf{t})=\mathbb{E}[\exp^{\mathfrak{i}\mathbf{t}^{T}Z}],\psi_{Y}(\mathbf{s})=\mathbb{E}[\exp^{\mathfrak{i}\mathbf{s}^{T}Y}],$ and $\psi_{Z,Y}(\mathbf{t},\mathbf{s})=\mathbb{E}[\exp^{\mathfrak{i}(\mathbf{t}^{T}Z+\mathbf{s}^{T}Y)}]$ be the characteristic functions of random vectors $Z\in\mathbb{R}^{d},Y\in\mathbb{R}^{m},$ and the pair $(Z,Y)$, respectively. The squared distance covariance $\mathcal{V}[Z,Y]$ is defined as
@@ 2.2.1 Distance covariance
where $c_{d}=\frac{\pi^{(d+1)/2}}{\Gamma((d+1)/2)}.$ Given $n$ i.i.d copies $\{Z_{i},Y_{i}\}_{i=1}^{n}$ of $(Z,Y)$, an unbiased estimator of $\mathcal{V}$ is the empirical distance covariance $\widehat{\mathcal{V}}_{n}$, which can be elegantly expressed as a $U$-statistic [45]
@@
$k\left(\left(\mathbf{z}_{1},\mathbf{y}_{1}\right),\ldots,\left(\mathbf{z}_{4},\mathbf{y}_{4}\right)\right)$ $=$ $\frac{1}{4}\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|\|\mathbf{y}_{i}-\mathbf{y}_{j}\|+\frac{1}{24}\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\left\|\mathbf{z}_{i}-\mathbf{z}_{j}\right\|\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{y}_{i}-\mathbf{y}_{j}\|$
@@
2.2.2 function class and norm-constrained neural network
Here, we assume that all optimal funcitons are $\beta$-Hölder smooth functions as stated in assumption 1 and 3 below. We aim to develop an approximation theory by utilizing the smoothness of optimal functions and obtain an explicit approximation error bound in terms of the network norm constraints.
@@
where $\partial^{\alpha}=\partial^{\alpha_{1}}\cdots\partial^{\alpha_{d}}$ with $\alpha=\left(\alpha_{1},\ldots,\alpha_{d}\right)^{\top}\in\mathbb{N}_{0}^{d}$ and $\|\alpha\|_{1}=\sum_{i=1}^{d}\alpha_{i}$. The Lipschitz class $\mathcal{H}^{1}:=\mathcal{H}([0,1]^{d},K_{2})$ is a special case of Hölder class with $\beta=1$. We adopt the Hölder class as the evaluation class, which encompasses the true functions. This is a common setup in nonparametric estimation [46].
Norm-constrained neural network. We consider a feedforward neural network (FNN) with a ReLU activation function to approximate the functions in the Hölder class. A neural network function $\phi:\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L+1}}$ is a function that can be parameterized as,
@@
where the activation function $\sigma(x):=x\lor 0$ is applied component-wisely and $T_{l}(\mathbf{X}):=A_{l}x+b_{l}$ is an affine transformation.
We first introduce a special class of neural network functions $\mathcal{SNN}(W,L)$ which contains functions of the form
$\hat{\phi}(\bm{x})=\hat{A}_{L}\sigma\left(\hat{A}_{L-1}\sigma\left(\cdots\sigma\left(\hat{A}_{0}\hat{\bm{x}}\right)\right)\right),\hat{\bm{x}}:=\left(\begin{array}[]{l}\bm{x}\\ 1\end{array}\right),$ (15)
where $\hat{A}_{\ell}\in\mathbb{R}^{\hat{N}_{\ell+1}\times\hat{N}_{\ell}}$ with $\hat{N_{0}}=d+1$ and $\max\left\{\hat{N}_{1},\ldots,\hat{N}_{L}\right\}=W$. Since these functions can also be written in the form (14) with $\bm{b}_{\ell}=\mathbf{0}$ for all $11\leq\ell\leq L-$ 1, it is obvious that $\mathcal{SNN}(W,L)\subseteq\mathcal{NN}(W,L)$. There is a natural way to introduce norm constraint on the weights: for any $K\geq 0$, we denote by $\mathcal{SN\mathcal{N}}(W,L,K)$ the set of functions in the form (15) that satisfies
$\prod_{\ell=0}^{L}\left\|\hat{A}_{\ell}\right\|\leq K.$
To introduce norm constraint for the class $\mathcal{NN}(W,L)$, we observe that any $\phi\in\mathcal{NN}(W,L)$ parameterized as (14) with $b_{L}=0$ can be written in the form
$\tilde{\phi}(\bm{x})=\tilde{A}_{L}\sigma\left(\tilde{A}_{L-1}\sigma\left(\cdots\sigma\left(\tilde{A}_{0}\tilde{\bm{x}}\right)\right)\right),\quad\tilde{\bm{x}}:=\left(\begin{array}[]{l}\bm{x}\\ 1\end{array}\right)$
with
$\tilde{A}_{L}=\left(A_{L},\mathbf{0}\right),\quad\tilde{A}_{\ell}=\left(\begin{array}[]{cc}A_{\ell}&\bm{b}_{\ell}\\ \mathbf{0}&1\end{array}\right),\ell=0,\ldots,L-1,$
and
$\prod_{\ell=0}^{L}\left\|\tilde{A}_{\ell}\right\|=\left\|A_{L}\right\|\prod_{\ell=0}^{L-1}\max\left\{\left\|\left(A_{\ell},\bm{b}_{\ell}\right)\right\|,1\right\}.$
Hence, we define the norm constrained neural network $\mathcal{NN}(W,L,K)$ as the set of functions $\phi_{\theta}\in\mathcal{NN}(W,L)$ of the form (14) that satisfies the following norm constraint on the weights
@@
We denote by $\mathcal{N}\mathcal{N}_{d_{1},d_{2}}(W,L)$ the set of ReLU FNN functions with width at most $W$, depth at most $L$ and its input (output) dimensions of $d_{1}$($d_{2}$). When the input dimension and output dimension are clear from contexts, we simply denote it by $\mathcal{N}\mathcal{N}(W,L)$. It is desirable to ensure bounded Lipschitz constant of neural networks for ease of analyse. It’s easy to see that any $f\in\mathcal{N}\mathcal{N}(W,L,K)$, $\|f\|_{L}<K$.
We set the network class used in Section 2.1.2-2.1.3 as follows: $\mathcal{H}=\mathcal{N}\mathcal{N}_{d,r}(W_{1},L_{1},K_{1})$, $\mathcal{G}=\mathcal{N}\mathcal{N}_{r,1}(W_{2},L_{2},K_{2})$ and $\mathcal{Q}=\mathcal{N}\mathcal{N}_{d^{*},1}(W_{3},L_{3},K)$.
@@ 3.1 Result for upstream training
(Regularity conditions on excess risk).
@@
The regularity conditions in Assumption 1 are mild. Firstly, we impose the boundedness assumptions on $\mathcal{X}$ and $\mathcal{Y}$ in (A1) for the sake of simplicity in presentation. This choice is facilitated by the ability to employ truncation techniques, given additional exponential decay tail assumptions on the distribution of the covariate and the noise, as elaborated in detail in [46]. Secondly, the assumption of a bounded target (A2) is standard in both parametric and nonparametric estimation. We required the network’s output to be bounded in (A3), which is achievable through a truncation implemented by a ReLU layer.
@@
Tripuraneni et al. [21, Theorem 4] presented a generalization bound for multitask transfer learning. However, their analysis is based on a well-specified neural network model, where the target function is also a network, without accounting for the approximation error as we have done.
@@ 3.2 Results for downstream prediction
(Regularity conditions for down-stream task). (A6) Each entries of $q^{*}$ is contained in $\mathcal{H}^{\beta}([-1,1]^{d},B)$. $\|\mathbf{F}_{T}^{*}\|_{\infty}\leq B$ and $\|\mathbf{A}^{*}\|_{F}\leq B$; (A7) $\ell(\cdot,y)$ is Lipschitz with respect to $y\in\mathcal{Y}$ and the Lipschitz constant is bounded by $L_{0}$, and $|\ell(\cdot,\cdot)|\leq B_{3}$; (A8) $\mathcal{W}_{1}(\mathbf{X},\mathbf{X}_{T})\leq\omega$ with $\omega\asymp\left(n^{-\frac{1}{2}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta+d+1}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)$; (A9) Consider $\mathcal{H}_{\eta}=\{h\in\mathcal{H}|\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}(X_{T}),\mathbf{h}^{*}(X_{T}))]\leq\eta,\eta\precsim\tilde{O}(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2})\},$ suppose $\sup_{\mathbf{h}\in\mathcal{H}_{\delta}}\mathbb{E}[\mathbb{E}_{Y_{T}}[\mathcal{W}_{1}(\mathbf{F}_{T}^{*}\mathbf{h}(X_{T})|Y_{T},\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})|Y_{T})]]\leq\nu\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}(X_{T}),\mathbf{h}^{*}(X_{T}))],$ where $\nu>0$ is a constant.
@@
Conditions (A6)-(A7) are standard requirements in the context of nonparametric supervised learning. Condition (A8) characterizes the similarity between the upstream and downstream tasks, a prerequisite for transfer feasibility. Let $\mathbb{P}_{\mathbf{X}}$ and $\mathbb{P}_{T}$ denote the marginal distributions of $\mathbf{X}$ and $X_{T}$ in the upstream and downstream tasks, respectively. In previous work, specifically in [21], the condition $\mathbb{P}_{\mathbf{X}}=\mathbb{P}_{T}$ or $\mathcal{W}_{1}(\mathbf{X},X_{T})=0$ was deemed necessary. In contrast, our assumption relaxes this requirement, allowing $\mathcal{W}_{1}(\mathbf{X},X_{T})$ to be small, indicating that a minor shift in input data is acceptable. Condition (A9) constitutes a technical requirement on local conditional distribution matching. Specifically, it implies that when the distribution of $\mathbf{h}(X_{T})$ closely approximates that of $\mathbf{h}^{*}(X_{T})$, the average conditional Wasserstein distance between the representations $\mathbf{h}(X_{T})$ and $\mathbf{h}^{*}(X_{T})$ is also small. In a recent study on domain adaptation, a similar, albeit more stringent, assumption known as conditional invariant component is employed, as outlined in [47].
@@
In the challenging scenario of “no transfer” where $\mathbf{F}_{T}^{*}=0$ and $d^{*}=d$, the convergence rate goes back to $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$. This rate aligns with the convergence rate achieved using ERM for learning the $d$-dimensional function $Q^{*}$ in classification [48] and is slower than the minimax rate observed in regression [49, 50, 51, 52, 53, 54, 55, 56]. Again, same with Remark 3 on Theorem 1, the convergence rate established in Theorem 3 is applicable to deep learning models characterized by over-parametrization. In the case of “partial transfer”, where $\mathbf{F}_{T}^{*}\neq 0$ and $Q^{*}(\cdot)=q^{*}(\mathbf{A}^{*}\cdot)$ with $d^{*}<d$, the convergence rate demonstrates improvement from $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$ to $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$, benefiting from the knowledge embedded in the upstream representation. In the case of “complete transfer” where $Q^{*}=0$, the convergence rate experiences further enhancement from $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$ to $\mathcal{O}(\frac{1}{\sqrt{m}})$. This enhancement is elaborated in detail in the following Theorem 4. Before this, we first give some assumptions.
@@
4 Experiment

### v3

@@ 2.1.2 Loss for representation learning
Notably, for any bijective mapping $\psi$ acting on $\mathbf{h}^{*}$, the composition $\psi\circ\mathbf{h}^{*}$ preserves the sufficiency and invariance properties required in (1)-(2). As a result, $\mathbf{h}^{*}(\mathbf{X})$ is not unique. To ensure uniqueness, following the standard approach in dimensionality reduction (Chen et al., 2024; Liu et al., 2024), we constrain the distribution of $\mathbf{h}^{*}(\mathbf{X})$ to match a given distribution $\mathcal{U}_{r}$. Specifically, we assume a uniform distribution for $\mathcal{U}_{r}$, as its existence can be guaranteed by optimal transport theory (Villani et al., 2009). Since the entries of $\mathcal{U}_{r}$ are independent, this promotes the disentanglement of $\mathbf{h}^{*}(\mathbf{X})$.
@@
$R(\mathbf{F},\mathbf{h})=$ $\mathbb{E}[\|Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S})\|^{2}]+\lambda\mathcal{V}[\mathbf{h}(\mathbf{X}),S]+\tau\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1},$ (10)
where the first term is the regression loss for (3), and $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ is a measure used to evaluate independence between $\mathbf{h}(\mathbf{X})$ and $S$ as required in (2), such as the distance covariance (detail definition see Section 2.2), and
@@
$\hat{R}(\mathbf{F},$ $\mathbf{h})=\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}_{i}\mathbf{h}(X_{i,j}))^{2}+\lambda\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]+\tau\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1},$ (11)
where $\sum_{j=1}^{p}n_{p}=n$, and $\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]$ is the empirical form of $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ (see Section 2.2 for detail), and
@@
denotes the empirical form of the Wasserstein distance, wherein $\{\xi_{i}\}_{i=1}^{n}$ denotes i.i.d random samples drawn from $\mathcal{U}_{r}$ and the function class $\mathcal{G}$ is chosen as the norm-constrained neural network class whose Lipschitz constants are bounded by $K_{3}$, see Section 2.2 for detail.
@@
$(\hat{\mathbf{F}},\hat{\mathbf{h}})=\operatorname*{arg\,min}_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\hat{R}(\mathbf{F},\mathbf{h}),$ (12)
@@ 2.1.3 Loss for downstream task
$(\hat{\mathbf{F}}_{T},\hat{Q})\in\operatorname*{arg\,min}_{(\mathbf{F}_{T},\mathbf{A},q)\in\mathcal{B}_{r}(B)\circ\mathcal{B}_{dd^{*}}(B)\circ\mathcal{Q}}\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ (13)
@@ 2.2 Function class and measure
Distance covariance. We recall the concept of distance covariance (Székely et al., 2007), which characterizes the dependence of two random variables. Let $\mathfrak{i}$ be the imaginary unit $(-1)^{1/2}$. For any $\mathbf{t}\in\mathbb{R}^{d}$ and $\mathbf{s}\in\mathbb{R}^{m}$, let $\psi_{Z}(\mathbf{t})=\mathbb{E}[\exp^{\mathfrak{i}\mathbf{t}^{T}Z}],\psi_{Y}(\mathbf{s})=\mathbb{E}[\exp^{\mathfrak{i}\mathbf{s}^{T}Y}],$ and $\psi_{Z,Y}(\mathbf{t},\mathbf{s})=\mathbb{E}[\exp^{\mathfrak{i}(\mathbf{t}^{T}Z+\mathbf{s}^{T}Y)}]$ be the characteristic functions of random vectors $Z\in\mathbb{R}^{d},Y\in\mathbb{R}^{m},$ and the pair $(Z,Y)$, respectively. The squared distance covariance $\mathcal{V}[Z,Y]$ is defined as
@@
where $c_{d}=\frac{\pi^{(d+1)/2}}{\Gamma((d+1)/2)}.$ Given $n$ i.i.d copies $\{Z_{i},Y_{i}\}_{i=1}^{n}$ of $(Z,Y)$, an unbiased estimator of $\mathcal{V}$ is the empirical distance covariance $\widehat{\mathcal{V}}_{n}$, which can be elegantly expressed as a $U$-statistic (Huo and Székely, 2016)
@@
$k\left(\left(\mathbf{z}_{1},\mathbf{y}_{1}\right),\ldots,\left(\mathbf{z}_{4},\mathbf{y}_{4}\right)\right)$ $=$ $\frac{1}{4}\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|\|\mathbf{y}_{i}-\mathbf{y}_{j}\|$
$+$ $\frac{1}{24}\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\left\|\mathbf{z}_{i}-\mathbf{z}_{j}\right\|\sum_{\begin{subarray}{c}1\leq i,j\leq 4\\ i\neq j\end{subarray}}\|\mathbf{y}_{i}-\mathbf{y}_{j}\|$
@@
where $\partial^{\alpha}=\partial^{\alpha_{1}}\cdots\partial^{\alpha_{d}}$ with $\alpha=\left(\alpha_{1},\ldots,\alpha_{d}\right)^{\top}\in\mathbb{N}_{0}^{d}$ and $\|\alpha\|_{1}=\sum_{i=1}^{d}\alpha_{i}$. The Lipschitz class $\mathcal{H}^{1}:=\mathcal{H}([0,1]^{d},K_{2})$ is a special case of Hölder class with $\beta=1$. We adopt the Hölder class as the evaluation class, which encompasses the true functions. This is a common setup in nonparametric estimation (Györfi et al., 2002).
Norm-constrained neural network. A neural network function $\phi:\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L+1}}$ is a function that can be parameterized as,
@@
where the activation function $\sigma(x):=x\lor 0$ is applied component-wisely and $T_{l}(\mathbf{X}):=A_{l}x+b_{l}$ is an affine transformation with $A_{l}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ and $b_{l}\in\mathbb{R}^{N_{l+1}}$ for $l=0,\dots,L$. The numbers $W=\max\{N_{1},\dots,N_{L}\}$ and $L$ are called the width and the depth of neural network, respectively. We denote by $\mathcal{N}\mathcal{N}_{d_{1},d_{2}}(W,L)$ the set of ReLU FNN functions with width at most $W$, depth at most $L$ and its input (output) dimensions of $d_{1}$($d_{2}$). When the input dimension and output dimension are clear from contexts, we simply denote it by $\mathcal{N}\mathcal{N}(W,L)$.
We define the norm constrained neural network $\mathcal{NN}(W,L,K)$ as the set of functions $\phi_{\theta}\in\mathcal{NN}(W,L)$ of the form (14) that satisfies the following norm constraint on the weights
@@
It’s easy to see that any $f\in\mathcal{N}\mathcal{N}(W,L,K)$, $\|f\|_{L}<K$. We set the network class used in Section  2.1.22.1.3 as follows: $\mathcal{H}=\mathcal{N}\mathcal{N}_{d,r}(W_{1},L_{1},K_{1})$, $\mathcal{G}=\mathcal{N}\mathcal{N}_{r,1}(W_{2},L_{2},K_{2})$ and $\mathcal{Q}=\mathcal{N}\mathcal{N}_{d^{*},1}(W_{3},L_{3},K)$.
@@ 3.1 Result for upstream training
(Regularity conditions on excess risk.)
@@
The regularity conditions in Assumption 1 are mild. Firstly, we impose the boundedness assumptions on $\mathcal{X}$ and $\mathcal{Y}$ in (A1) for the sake of simplicity in presentation. This choice is facilitated by the ability to employ truncation techniques, given additional exponential decay tail assumptions on the distribution of the covariate and the noise, as elaborated in detail in Györfi et al. (2002). Secondly, the assumption of a bounded target (A2) is standard in both parametric and nonparametric estimation. We required the network’s output to be bounded in (A3), which is achievable through a truncation implemented by a ReLU layer.
@@
Tripuraneni et al. (2020, Theorem 4) presented a generalization bound for multitask transfer learning. However, their analysis is based on a well-specified neural network model, where the target function is also a network, without accounting for the approximation error as we have done.
@@ 3.2 Results for downstream prediction
(Regularity conditions for down-stream task). (A6) Each entries of $q^{*}$ is contained in $\mathcal{H}^{\beta}([-1,1]^{d},B)$. $\|\mathbf{F}_{T}^{*}\|_{\infty}\leq B$ and $\|\mathbf{A}^{*}\|_{F}\leq B$; (A7) $\ell(\cdot,y)$ is Lipschitz with respect to $y\in\mathcal{Y}$ and the Lipschitz constant is bounded by $L_{0}$, and $|\ell(\cdot,\cdot)|\leq B_{3}$; (A8) $\mathcal{W}_{1}(\mathbf{X},\mathbf{X}_{T})\leq\omega$ with $\omega\asymp\left(n^{-\frac{1}{2}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta+d+1}{2d+3\beta}}\mathbb{I}_{\beta>2}\right)$; (A9) Consider $\mathcal{H}_{\eta}=\{h\in\mathcal{H}|\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}(X_{T}),\mathbf{h}^{*}(X_{T}))]\leq\eta,\eta\precsim\tilde{O}(n^{-\frac{\beta}{2(d+1+\beta)}}\mathbb{I}_{\beta\leq 2}+n^{-\frac{\beta}{2d+3\beta}}\mathbb{I}_{\beta>2})\}.$
Suppose $\sup_{\mathbf{h}\in\mathcal{H}_{\delta}}\mathbb{E}[\mathbb{E}_{Y_{T}}[\mathcal{W}_{1}(\mathbf{F}_{T}^{*}\mathbf{h}(X_{T})|Y_{T},\mathbf{F}_{T}^{*}\mathbf{h}^{*}(X_{T})|Y_{T})]]\leq\nu\mathbb{E}[\mathcal{W}_{1}(\mathbf{h}(X_{T}),\mathbf{h}^{*}(X_{T}))],$ where $\nu>0$ is a constant.
@@
Conditions (A6)-(A7) are standard requirements in the context of nonparametric supervised learning. Condition (A8) characterizes the similarity between the upstream and downstream tasks, a prerequisite for transfer feasibility. Let $\mathbb{P}_{\mathbf{X}}$ and $\mathbb{P}_{T}$ denote the marginal distributions of $\mathbf{X}$ and $X_{T}$ in the upstream and downstream tasks, respectively. In previous work, specifically in Tripuraneni et al. (2020), the condition $\mathbb{P}_{\mathbf{X}}=\mathbb{P}_{T}$ or $\mathcal{W}_{1}(\mathbf{X},X_{T})=0$ was deemed necessary. In contrast, our assumption relaxes this requirement, allowing $\mathcal{W}_{1}(\mathbf{X},X_{T})$ to be small, indicating that a minor shift in input data is acceptable. Condition (A9) constitutes a technical requirement on local conditional distribution matching. Specifically, it implies that when the distribution of $\mathbf{h}(X_{T})$ closely approximates that of $\mathbf{h}^{*}(X_{T})$, the average conditional Wasserstein distance between the representations $\mathbf{h}(X_{T})$ and $\mathbf{h}^{*}(X_{T})$ is also small. In a recent study on domain adaptation, a similar, albeit more stringent, assumption known as conditional invariant component is employed, as outlined in Wu et al. (2023).
@@
In the challenging scenario of “no transfer” where $\mathbf{F}_{T}^{*}=0$ and $d^{*}=d$, the convergence rate goes back to $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$. This rate aligns with the convergence rate achieved using ERM for learning the $d$-dimensional function $Q^{*}$ in classification (Shen et al., 2022) and is slower than the minimax rate observed in regression (Schmidt-Hieber, 2019; Nakada and Imaizumi, 2020; Farrell et al., 2021; Chen et al., 2022; Kohler et al., 2022; Fan and Gu, 2023; Bhattacharya et al., 2023; Jiao, Shen, Lin and Huang, 2023). Again, same with Remark 3 on Theorem 1, the convergence rate established in Theorem 3 is applicable to deep learning models characterized by over-parametrization. In the case of “partial transfer”, where $\mathbf{F}_{T}^{*}\neq 0$ and $Q^{*}(\cdot)=q^{*}(\mathbf{A}^{*}\cdot)$ with $d^{*}<d$, the convergence rate demonstrates improvement from $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$ to $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$, benefiting from the knowledge embedded in the upstream representation. In the case of “complete transfer” where $Q^{*}=0$, the convergence rate experiences further enhancement from $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$ to $\mathcal{O}(\frac{1}{\sqrt{m}})$. This enhancement is elaborated in detail in the following Theorem 4. Before this, we first give some assumptions.
@@
4 Experiment on four real classification datasets and one regression dataset